# Fully stochastic trust-region methods with Barzilai-Borwein steplengths

### v1

@@ 4.3.2 Datasets MNIST and CIFAR10
We present the averaged accuracies obtained for each of the five epochs performed in the experiments and for each triplet $(\alpha,\gamma_{1},\gamma_{2})=(1,\gamma_{1},\gamma_{2})$. As shown in Table 4, we first observe that the variants of TRishBB employed the unconstrained trust-region solution less frequently than for datasets (a1a, w1a, cina) but the percentage of unconstrained steps increases with the value of $\alpha$.
@@
Figure 11Figure 13 display the averaged accuracy obtained with the dataset MNIST. TRishBB_v2 compares well with TRishBB_v1 even though it employed a smaller percentage of unconstrained steps. Both TRishBB_v1 and TRishBB_v2 are quite insensitive to the values of $\gamma_{1}$ and $\gamma_{2}$. TRishBB_v3 selected the unconstrained steps less than $1\%$ of the iterations and its performance is similar to the performance of TRish.
@@
Figure 14Figure 16 show the averaged accuracies with the dataset CIFAR10. The main observation is that the proposed TRishBB variants are less sensitive than TRish to the choice of $\gamma_{1}$ and $\gamma_{2}$, see e.g., experiments Ex432 and Ex433. Except for the first epoch, TRishBB_v2 performs better than TRish and than the other two variants. Algorithm TRishBB_v1 generally performed better than Algorithm TRishBB_v3 but it is not a clear winner against TRish. TRishBB_v3 is again comparable to TRish since the percentage of unconstrained solution taken is less than $2\%$.
Figure 17 shows that the rule for generating $\mu_{k}$ in TRishBB_v2 yields small values. The values of $\mu_{k}$ generated by TRishBB_v1 are smaller than in TRishBB_v2 , in accordance with the high number of unconstrained steps taken by the two algorithms, but the overall results indicate that the rule for building the steplength $\mu_{k}$ in TRishBB_v2 is the most effective. The values of $\mu_{k}$ built in TRishBB_v3 are the largest at all iterations.
@@ 5 Conclusion
We introduced the Algorithm TRishBB applicable to finite-sum minimization problems. It is a stochastic trust-region method that combines the TRish methodology with stochastic Barzilai-Borwein steplengths to inject cheap second-order information. We presented theoretical analysis of TRishBB and devised three practical variants differing in the computation of the BB steplength. TRishBB_v1 mimics the implementation of a standard BB rule, TRishBB_v2 employs an accumulated BB parameter per iteration, TRishBB_v3 retrieves the second-order information by the empirical Fisher Information matrix.
Numerical results demonstrated that TRishBB_v2 is a reliable algorithm that outperforms the other variants and that it is quite insensitive to the choice of the hyperparameters $\alpha$, $\gamma_{1}$ and $\gamma_{2}$; also for large value of the parameter $\alpha$, TRishBB_v2 performs better than the first-order TRish method.
@@
Figure 11: MNIST: Testing accuracy of TRishBB_v1 with $(\alpha,\gamma_{1},\gamma_{2})=(1,\gamma_{1},\gamma_{2})$, $m=20$, and $|\mathcal{N}_{k}|=128$
Figure 12: MNIST: Testing accuracy of TRishBB_v2 with $(\alpha,\gamma_{1},\gamma_{2})=(1,\gamma_{1},\gamma_{2})$, $m=N_{b}$, and $|\mathcal{N}_{k}|=128$ using damping relaxed combination
Figure 13: MNIST: Testing accuracy of TRishBB_v3 with $(\alpha,\gamma_{1},\gamma_{2})=(1,\gamma_{1},\gamma_{2})$, $m=5$, and $|\mathcal{N}_{k}|=128$
Figure 14: CIFAR10: Testing accuracy of TRishBB_v1 with $(\alpha,\gamma_{1},\gamma_{2})=(1,\gamma_{1},\gamma_{2})$, $m=20$, and $|\mathcal{N}_{k}|=128$
Figure 15: CIFAR10: Testing accuracy of TRishBB_v2 with $(\alpha,\gamma_{1},\gamma_{2})=(1,\gamma_{1},\gamma_{2})$, $m=N_{b}$, and $|\mathcal{N}_{k}|=128$ using damping relaxed combination
Figure 16: CIFAR10: Testing accuracy of TRishBB_v3 with $(\alpha,\gamma_{1},\gamma_{2})=(1,\gamma_{1},\gamma_{2})$, $m=5$, and $|\mathcal{N}_{k}|=128$
Figure 17: Average values of $\mu_{k}$ with triplets $(\alpha,\gamma_{1},\gamma_{2})=(1,\gamma_{1},\gamma_{2})$ along iterations for MNIST and CIFAR10.
@@ A Some intermediate results
B Proofs from §Section 3

### v2

@@ 4.3.2 Datasets MNIST and CIFAR10
We present the averaged accuracies obtained for each of the five epochs performed in the experiments and for each triplet $(\alpha,\gamma_{1},\gamma_{2})=(1,\gamma_{1},\gamma_{2})$. As shown in Table 2, the variants of TRishBB employed the unconstrained trust-region solution less frequently than for datasets a1a, w1a and cina, but the percentage of unconstrained steps increases with the value of $\alpha$.
@@
Figure 5 displays the averaged accuracy obtained with the dataset MNIST. TRishBB_v2 compares well with TRishBB_v1 even though it employed a smaller percentage of unconstrained steps. Both TRishBB_v1 and TRishBB_v2 are considerably less sensitive to the values of $\gamma_{1}$ and $\gamma_{2}$ compared to TRish. TRishBB_v3 exhibits performance similar to that of TRish, which is not surprising, as it selects the unconstrained step in only approximately 15.23% of the iterations. Nevertheless, TRishBB_v3 outperforms TRish at the first epoch due to the predominance of BB steplengths in that phase; specifically, Table 4 shows the number of BB steplengths taken at the end of the first epoch compared to the total number taken at the end of the fifth epoch.
@@
Figure 6 shows the averaged accuracies with the dataset CIFAR10. The main observation is that the proposed TRishBB variants are less sensitive than TRish to the choice of $\gamma_{1}$ and $\gamma_{2}$, see e.g., experiments Ex431, Ex432 and Ex433. Since accuracies of TRishBB variants and TRish are increasing during the epochs, we focus on the fifth epoch and observe that TRishBB_v2 performs better than all other algorithms. TRishBB_v1 and TRishBB_v3 performed better than TRish limited to some choices of $\gamma_{1}$ and $\gamma_{2}$.
Figure 4 shows that the values of $\mu_{k}$ generated by TRishBB_v1 are smaller than TRishBB_v2 and TRishBB_v3, consistently with Table 2. Although TRishBB_v1 takes many BB steplengths, TRishBB_v2 stands out as the best-performing variant as in the previous set of experiments.
Table 4: The number of BB steplengths $\mu_{k}$ taken by TRishBB_v3 on MNIST at the end of first epoch vs. the total number taken at the end of the fifth epoch.
Experiments E411 E412 E413 E421 E422 E423 E431 E432 E433
[…]
Figure 5: MNIST: Testing accuracy, $(\alpha,\gamma_{1},\gamma_{2})=(1,\gamma_{1},\gamma_{2})$, $|\mathcal{N}_{k}|=128$. Top: TRishBB_v1, $m=20$. Middle: TRishBB_v2, $m=468$. Bottom: TRishBB_v3, $m=468$
Figure 6: CIFAR10: Testing accuracy, $(\alpha,\gamma_{1},\gamma_{2})=(1,\gamma_{1},\gamma_{2})$, $|\mathcal{N}_{k}|=128$. Top: TRishBB_v1, $m=20$. Middle: TRishBB_v2, $m=390$. Bottom: TRishBB_v3, $m=390$.
4.3.3 Average testing loss obtained by TRishBB_v2
Sections 4.3.1 and 4.3.2 indicate that TRishBB_v2 is the overall best variant of TRishBB. In this section, we present the average testing loss vs. the number of gradient evaluations obtained with both TRishBB_v2 and TRish over five epochs. We consider specific configurations ExIJK for the triplet $(\alpha,\gamma_{1},\gamma_{2})$ described in Section 4.2. The results shown in Figure 7 refer to the highest average testing accuracy of these algorithms in Table 3 for the largest value of $\alpha$. Specifically, TRishBB_v2 achieves its best performance with configuration Ex511 on a1a, w1a, and cina, Ex422 on MNIST, and Ex413 on CIFAR10, while TRish attains its highest accuracy with Ex532 on a1a, Ex524 on w1a, Ex531 on cina, Ex411 on MNIST, and Ex421 on CIFAR10. The figure shows that our method reaches lower testing loss values earlier than TRish and that in the solution of a1a, w1a, and cina the value of the testing loss at termination is largely different and in favour of TRishBB_v2.
Figure 7: Average testing loss of TRishBB_v2 vs. number of gradient evaluations (NGE) compared to TRish across five datasets.
4.3.4 Comparison to other algorithms
As discussed in Sections 2.2 and 2.3, TRishBB_v2 borrows features from SGD-BB method given in [11], while TRishBB_v3 borrows ideas from AdaQN Algorithm given in [34]. In this section, we compare the performance of these algorithms. For TRishBB_v2 and TRishBB_v3, all hyper-parameters are set as described in Section 4.2.
TRishBB_v2 vs. SGD-BB
We implemented SGD-BB based on Algorithm 3 in [11] allowing mini-batch size greater than one. We set the cardinality of $\mathcal{N}_{k}$, the scalar $\beta$ and the cyclic update parameter $m$ as described in Section 4.2 for the experiments with TRishBB_v2.
We remark that the original SGD-BB algorithm did not apply thresholding for $\mu_{k}$ but our experiments demonstrated that thresholding is necessary for robustness, along with fine tuning for the initial value $\mu_{0}$. Specifically, initially, we ran SGD-BB with $\mu_{0}=1$ which was the setting used in TRishBB_v2 but the BB steplength generated in SGD-BB were too large causing failures in the fist epoch. To mitigate this issue, we ran the algorithm with $\mu_{0}=10^{-2}$ for a1a, w1a, and cina, and with $\mu_{0}=10^{-3}$ for MNIST and CIFAR10. Nonetheless, these smaller values for $\mu_{0}$ were still insufficient to provide good results in the next epochs. Hence, we applied the thresholding of $\mu_{k}$ as in TRishBB_v2 and verified that setting $\mu_{\min}=10^{-5}$ and $\mu_{\max}=10^{-1}$ provided convergence for all datasets.
Table 5 compares the performance of SGD-BB against TRishBB_v2 performed with the largest values of $\alpha$ indicated in Sections 4.3.1 and 4.3.2. The highest average testing accuracy achieved by TRishBB_v2 per epoch is close to or better than the accuracy achieved by tuned SGD-BB. Indeed, in SGD-BB a careful adjustment of both the initial stepsize $\mu_{0}$ and the threshold $\mu_{\max}$ was necessary to obtain good performance and the average testing accuracy may differ greatly from one epoch to the other.
Table 5: Highest average testing accuracy per epoch obtained by TRishBB_v2 with the largest $\alpha$ and by SGD-BB.
[…]
TRishBB_v2 83.14 83.52 79.71 80.94 80.77
SGD-BB 82.63 81.73 60.03 61.76 54.57
[…]
TRishBB_v2 89.38 89.44 89.47 89.53 89.57
SGD-BB 89.38 89.46 89.54 89.45 87.74
[…]
TRishBB_v2 90.07 90.99 91.35 91.45 91.20
SGD-BB 89.52 91.01 91.32 83.99 49.21
[…]
TRishBB_v2 98.30 98.61 98.73 98.71 98.81
SGD-BB 93.44 95.01 73.08 98.26 98.70
[…]
TRishBB_v2 48.18 58.42 62.52 65.01 65.98
SGD-BB 20.54 27.49 19.57 51.67 57.83
TRishBB_v3 vs. AdaQN
Our experiments with AdaQN were performed using adaQN.m from the package minSQN, available at https://github.com/keskarnitish/minSQN/blob/master/%2Bmethods/adaQN.m. As suggested therein and in the reference paper [34], we set the hyper-parameters
We set the limited-memory size of the L-BFGS Hessian approximations $H_{k}$ to 202020, and the Fisher information storage size $m_{F}$ to 100100100. AdaQN provides two strategies for initializing $H_{0}$. We adopted the default option, RMSprop-like initialization, which showed better performance in our experiments compared to the Adagrad-like strategy.
The cyclic iteration parameter $L$ in AdaQN, corresponding to our parameter $m$ in TRishBB_v3, needs fine tuning and we observed that it is highly sensitive to the choice of random seed during hyper-parameter tuning. We fixed its value in the set $L\in\{2,5,10,20,64,m\}$ where the parameter $m$ is the same as in TRishBB_v3. The remaining hyper-parameter values were adopted from [34] and the implementation mentioned above. Following this implementation, the stepsize $\alpha$ in AdaQN was sampled from a log-uniform distribution over the interval $[10^{-6},10^{2}]$. We tuned AdaQN using 20 different random seeds for the values of $L$ and $\alpha$, as part of the standard procedure in the adaQN.m routine, with each run lasting 5 epochs. Using the resulting optimal hyper-parameters $L$ and $\alpha$, we obtained the results presented in Table 6 where AdaQN is compared with TRishBB_v3 using the largest $\alpha$ indicated in Sections 4.3.1 and 4.3.2.
Table 6: Highest average testing accuracy per epoch obtained by TRishBB_v3 with the largest $\alpha$ and by AdaQN with its tuned $(L,\alpha)$.
[…]
TRishBB_v3 83.19 83.40 82.60 78.65 79.42
AdaQN $(L=10,\,\alpha=1.13\times 10^{-3})$ 76.62 78.83 80.34 80.52 80.48
[…]
TRishBB_v3 89.34 89.42 89.51 89.47 89.24
AdaQN ($L=20,\,\alpha=8.41\times 10^{-3}$) 48.27 47.78 47.65 47.70 47.57
[…]
TRishBB_v3 90.00 91.18 90.07 88.58 85.05
AdaQN ($L=5,\,\alpha=7.21\times 10^{-3}$) 88.73 87.76 87.92 88.25 88.14
[…]
TRishBB_v3 98.28 97.91 98.26 98.25 98.59
AdaQN ($L=468,\,\alpha=6.50\times 10^{-4}$) 98.13 98.34 98.66 98.82 98.86
[…]
TRishBB_v3 49.55 58.45 61.78 63.80 64.89
AdaQN ($L=64,\,\alpha=4.73\times 10^{-4}$) 54.65 60.65 62.99 66.02 66.46
The table shows that the highest average testing accuracy achieved by TRishBB_v3 per epoch over a1a, w1a, and cina is generally larger than the accuracy achieved by AdaQN, in particular for earlier epochs. On the other hand, the performance of both TRishBB_v3 and AdaQN is close to each other for MNIST while AdaQN outperforms TRishBB_v3 in the solution of CIFAR10. In MNIST and CIFAR10 problems, TRishBB_v3 could meet rarely the iterations at which the BB steplength is computed, see Table 2; in addition, the good performance of AdaQN is motivated by the use of an L-BFGS-based Hessian approximation coupled with a step acceptance criterion based on sampled function evaluations.
4.3.5 The cyclic iteration parameter $m$ in TRishBB_v3
We conclude our experimental analysis, focusing on the cyclic iteration parameter $m$ in TRishBB_v3. In all experiments conducted we have fixed $m=N_{b}$ with $N_{b}=\lfloor N/|\mathcal{N}_{k}|\rfloor$ ($N_{b}$ equal to 25, 38, 156, 468, and 390 for a1a, w1a, cina, MNIST, and CIFAR10, respectively). To assess the sensitivity of TRishBB_v3 to the hyper-parameter $m$, we report in Figures 89 additional results obtained with $m\in\{5,20,64,N_{b}\}$, i.e., the values used for tuning Ada-QN; to improve readability in the figures we have eliminated the results corresponding to $m=2$ and $m=10$ because $m=5$ sufficiently represents the impact of small values of $m$. These figures show that for large values of $m$, $m=64$ and $m=N_{b}$, the performance is not highly sensitive to the choice of $\gamma_{1}$ and $\gamma_{2}$ and that the setting $m=N_{b}$ provides high accuracy at some epochs for all tests, making unnecessary to tune a further hyper-parameter.
@@ 5 Conclusion
We introduced the Algorithm TRishBB applicable to finite-sum minimization problems. It is a stochastic trust-region method that combines the TRish methodology with stochastic Barzilai-Borwein steplengths to inject cheap second-order information. We presented theoretical analysis of TRishBB and devised three practical variants differing in the computation of the BB steplength. TRishBB_v1 mimics the implementation of a standard BB rule, TRishBB_v2 employs an accumulated BB parameter per iteration, TRishBB_v3 retrieves the second-order information by the accumulated Fisher Information matrix. Numerical results demonstrated that, for large values of $\alpha$, TRishBB_v2 and TRishBB_v3 are quite insensitive to the hyper-parameters and performs better than the first-order TRish method. We also observe that they compare well with two stochastic quasi-Newton methods from the literature. Taking into account that TRishBB_v3 is applicable as long as it is feasible the use of the empirical Fischer Information, TRishBB_v2 appears to be a robust and general algorithm.
@@
Figure 8: Testing average accuracy of TRishBB_v3 with $(\alpha,\gamma_{1},\gamma_{2})=(10,\gamma_{1},\gamma_{2})$, $m\in\{5,20,64,N_{b}\}$. Top: a1a, $N_{b}=25$. Middle: w1a, $N_{b}=38$. Bottom: cina, $N_{b}=156$.
Figure 9: Testing average accuracy of TRishBB_v3 with $(\alpha,\gamma_{1},\gamma_{2})=(1,\gamma_{1},\gamma_{2})$, $m\in\{5,20,64,N_{b}\}$. Top: MNSIT, $N_{b}=468$. Bottom: CIFAR10, $N_{b}=390$.
@@ A Some intermediate results
B Proofs from Section 3
@@
C Classification Problems
Table 7: Classification test problems.
Dataset $\mathbf{(N,{N}_{t})}$ $\mathbf{d}$ $\mathbf{n}$ Type $\mathbf{f_{i}(x)}$ NN
a1a [36] (1605,29351)(1605,29351)( 1605 , 29351 ) 123123123 123123123 Binary Logistic regression -
w1a [36] (2477,47272)(2477,47272)( 2477 , 47272 ) 300300300 300300300 Binary Logistic regression -
cina [36] (10000,6033)(10000,6033)( 10000 , 6033 ) 132132132 132132132 Binary Logistic regression -
MNIST [37] (60000,10000)(60000,10000)( 60000 , 10000 ) 28×28×128$\times$ 28$\times$ 128 × 28 × 1 431030431030431030 Multi Cross-entropy NET-1
CIFAR10 [38] (50000,10000)(50000,10000)( 50000 , 10000 ) 32×32×332$\times$ 32$\times$ 332 × 32 × 3 545098545098545098 Multi Cross-entropy NET-2
Table 7 summarizes the five datasets used in our experiments for both binary and multi-class classification tasks in Section 4.1. Each test dataset with $C$ possible classes is divided into training and testing sets, denoted as $\{(a_{i},b_{i})\,|i\in\mathcal{N}\}$ and $\{(a_{i},b_{i})\,|i\in\mathcal{N}_{t}\}$, with $|\mathcal{N}|=N$ and $|\mathcal{N}_{t}|={N}_{t}$ samples, respectively; the values of $N$ and ${N}_{t}$ are reported in the table. The vector $a_{i}\in\mathbb{R}^{d}$ represents the $i$-th data with $d$ features while $b_{i}$ is the corresponding categorical label associated.
The dataset cina was normalized such that each feature vector has components in the interval [0,1][0,1][ 0 , 1 ]. For MNIST and CIFAR10, besides the zero-one rescaling, we implemented z-score normalization to have zero mean and unit variance. Precisely, all $N$ training images, (3-D arrays of size $\sigma_{1}\times\sigma_{2}\times\sigma_{2}$ where $\sigma_{1}$, $\sigma_{2}$ and $\sigma_{3}$ denote the height, width, and number of channels of the images, respectively), undergo normalization by subtracting the mean (a 3-D array) and dividing by the standard deviation (a 3-D array) component-wise. Data in the testing set are also normalized using the same statistics as in the training data. The architecture of the neural network (NN) is described in Table 8.111The compound $Conv(5,20,1,0){/}ReLU{/}MaxPool(2,2,0)$ indicates a convolutional layer ($Conv$) using 202020 filters of size $\times 5$ 5, stride 111, and no padding, a nonlinear activation function ($ReLu$) and, a 2-D max-pooling layer with a channel of size $\times 2$ 2, stride 2 and no padding. The syntax $FC(C){/}Softmax$ denotes a fully connected layer of $C$ neurons followed by the $softmax$ layer. For MNIST dataset with grayscaled images in $C=10$ classes (digits 0–9), and CIFAR10 with RGB images in $C=10$ categories (e.g., dogs and ships), we used NET-1 and NET-2 respectively.
Table 8: NN architecture.
NET-1 $Input(d)$
$Conv(5,20,1,0){/}ReLU{/}MaxPool(2\times 2,2,0)$
$Conv(5,50,1,0){/}ReLU{/}MaxPool(2\times 2,2,0)$
$FC(500){/}ReLU$
$FC(C){/}Softmax$
NET-2 $Input(d)$
$Conv(3,32,1,same){/}ReLU{/}AveragePool(2\times 2,2,0)$
$Conv(3,64,1,same){/}ReLU{/}AveragePool(2\times 2,2,0)$
$FC(128){/}ReLU$
$FC(C){/}Softmax$